# Characterization of the western Pictor A hotspot in the hard X-rays with NuSTAR

### v1

@@ Abstract
The origin of X-ray emission from the resolved kiloparsec-scale jets of many active galactic nuclei (AGN) remains uncertain, particularly where the X-ray emission is separate from the radio-optical synchrotron component. Possible explanations include inverse-Compton upscattering of the Cosmic Microwave Background (IC-CMB) and synchrotron emission from a second electron population, alternatives which imply very different physical conditions within the jet. Until now, X-ray studies of resolved jets have been restricted to the soft X-rays (below 8 keV), often showing a hard spectral index indicating a spectral peak at higher energies. Here we present NuSTAR and Chandra observations of the nearby powerful radio galaxy Pictor A, in which we clearly detect the western hot spot at 4’ from the host galaxy, the first detection of a kpc-scale jet above 10 keV. The NuSTAR 320 keV spectrum is best fit by a powerlaw of index . Using a maximum likelihood model, we confirm previous reports of variations in the soft X-ray flux detected by Chandra over the 2000 to 2015 period, at a 6.7 level, higher than the 3.4 significance level. This rises to in the common 3-8 keV band using the full 22-year span of Chandra and NuSTAR observations. The variability of Pictor A rules out IC-CMB as the source of the X-ray emission, and the NuSTAR observations are in keeping with a synchrotron origin for the X-ray emission from the western hotspot of Pictor A, with no sign of a cutoff to the flat spectrum.
@@ 1 Introduction
In addition to (possible) terminal hot spots, AGN jets also generally feature discrete emitting regions called knots. Both the knots and hot spots are primarily detected at radio and sometimes IR/optical frequencies where they emit synchrotron radiation from electrons gyrating within the jet magnetic field. For this reason, jetted AGN are more ‘radio-loud’ compared to the majority of non-jetted AGN. Sources where the jet is aligned in the plane of the sky are known as radio galaxies. Radio galaxies are often divided into two classifications based on jet morphology: Fanaroff Riley type I (FR I; Fanaroff & Riley, 1974) jets tend to be shorter, less collimated, and brightest near the central engine with the jet typically terminating in large, darkened plumes of plasma, while FRII jets are highly collimated and brightest at their terminating hot spots.
@@
One of the major discoveries of Chandra was the detection of ‘anomalously’ bright X-ray emission from resolved radio jets on kpc scales (Chartas et al., 2000; Sambruna et al., 2001). In some FRI jets such as M87 (Harris et al., 2003; Sun et al., 2018) and Cen A (Kraft et al., 2002; Snios et al., 2019), the X-ray emission is roughly consistent with a single spectral component extending from radio to optical to X-rays. In other, often more powerful jets (Harris & Krawczynski, 2006, 2007; Harris et al., 2010; Marshall et al., 2010) the X-ray emission is often ‘anomalous’, meaning that the X-ray flux is too bright and hard-spectrum to be an extension of the primary radio-optical synchrotron component.
While it was early on concluded that the X-ray emission expected due to synchrotron self-Compton emission is far below the observed X-ray flux (e.g., Schwartz et al., 2000), Tavecchio et al. (2000) and Celotti et al. (2001) independently showed that relativistically boosted inverse-Compton upscattering of cosmic microwave background photons (IC-CMB) could produce sufficient X-rays, provided that the kpc-scale jet is closely aligned to the line-of-sight () and remains highly relativistic (bulk Lorentz factor ). In this case, IC-CMB emission can be Doppler boosted to levels in agreement with observation.
Though often adopted as the default mechanism for high X-ray fluxes in kpc-scale jets, there is now significant evidence in the literature suggesting that IC-CMB is, at best, only capable of explaining the emission for a subset of sources. Observations incompatible with the model include high optical/UV polarization fractions in the second spectral component (Cara et al., 2013), spatial offsets between X-ray and radio emitting regions in many sources (Sambruna et al., 2004; Jester et al., 2006; Kataoka et al., 2008; Reddy et al., 2021a), and inconsistent values of and as measured from jet-counterjet flux ratios (Kataoka et al., 2008; Clautice et al., 2016; Hardcastle et al., 2016) as well as a lack of proper motions in the kpc-scale jet (Meyer et al., 2016). The IC-CMB model also predicts a significant steady GeV -ray flux (Georganopoulos et al., 2006), at a level violated by deep gamma-ray band upper limits from the Fermi/LAT for more than a dozen sources (Meyer et al., 2015, 2017; Breiding et al., 2017, 2023). Despite these challenges, IC-CMB is still frequently adapted as the default model (e.g. Kharb et al., 2012; Gentry et al., 2015). It may, in fact, be the correct explanation for anomalous X-ray emission in sources at high redshift (), where the CMB energy density is higher (e.g. Siemiginowska et al., 2003; Hodges-Kluck et al., 2021), and rare low redshift sources which satisfy the aforementioned requirements for and (e.g. Meyer et al., 2019).
The main alternative explanation is that the second, X-ray-producing component is synchrotron radiation from a second, high-energy electron population, distinct from the radio-emitting population (Hardcastle et al., 2004; Jester et al., 2006). This would require in situ particle acceleration to accelerate the electrons within the X-ray bright knots and hot spots to the necessary TeV energies as the radiating particles have cooling timescales much smaller than the transport time down the jet.
The source of the kpc-scale X-ray emission for powerful jets is a longstanding open question with the competing IC-CMB and synchrotron models often implying orders-of-magnitude differences in jet kinetic power and magnetic field strength. Determining to what extent each mechanism contributes across the X-ray jet population is thus important also for understanding the energy content of jets and their impact.
@@
Given the unsolved origin of the X-ray emission, it is useful to attempt to extend the spectral coverage with observations in the hard X-ray range ( keV). If the emission is synchrotron in origin, locating the turnover point where the spectrum begins to fall can place constraints on the maximum electron energy. This energy range is covered by NuSTAR, but observations of kpc-scale jets are complicated by the fact that NuSTAR lacks the angular resolution (FWHM7.5”) to easily distinguish the kpc-scale jet from the core in the case of most sources. Given these limitations, the outer jet and western hot spot of FR II radio galaxy Pictor A (; 0.69 kpc/”) is nearly the only jet with the sufficient angular separation () from the core to be a viable target for hard X-ray study and the only one bright enough to be detected by NuSTAR in less than approximately 1 Ms of exposure.
@@
Pictor A is among the most well-studied X-ray jets in the literature, and has been the subject of a number of detailed studies over the past two decades (e.g., Wilson et al., 2001; Hardcastle et al., 2004; Hardcastle & Croston, 2005; Migliori et al., 2007; Tingay et al., 2008; Marshall et al., 2010; Hardcastle et al., 2016; Thimmappa et al., 2020). The X-ray jet was first reported by Wilson et al. (2001) and was noted to be X-ray-anomalous by Hardcastle et al. (2004). Despite its bright X-ray emission, it is not a source which we would expect a significant IC-CMB component as it has a low redshift, its (approaching) western jet is not closely inclined to our line-of-sight (), and its counterjet is clearly visible (Hardcastle & Croston, 2005). A beamed IC-CMB emission origin for the X-rays from the jet or hot spot would require magnetic field strengths significantly below equipartition values (Hardcastle et al., 2004). The deepest observation of the jet and hot spot to date in the soft X-rays combined 464 ks of Chandra exposure on the western hot spot, finding a spectrum near-equally consistent with a single flat powerlaw spectrum with or a steepening, broken powerlaw spectrum with , , and a break energy at around 2 keV (Hardcastle et al., 2016). One of the motivations for our deep NuSTAR observation was to clarify the presence and location of a possible spectral break in the broad X-ray band.
@@
The jet of Pic A is also noteworthy for the detection of variability in the X-rays. In 2010, Marshall et al. (2010) analyzed three Chandra observations taken in 2000, 2002, and 2009 of 26, 96, and 54 ks of exposure time respectively. They found that the soft X-ray flux of two features located at 34 and 49 kpc away from the core decreased significantly ( and respectively) over the course of the observations. In the later Hardcastle et al. (2016) study (with an additional 6 Chandra observations) found the hot spot X-ray flux decreased by over a time period of approximately one month, and confirmed 3 variability in the jet. Very recently, Pictor A was included in a large archival study of Chandra jet observations where at least five regions in the jet, including the hotspot, were found to be variable (Meyer et al., 2023); we further expand on the results of that study here using the same methods.
The IC-CMB mechanism requires low-energy electrons () with cooling timescales on the order of yrs (Worrall, 2009), so the observed variability in the X-ray flux is clearly incompatible with such an origin.
@@
In this paper, we report the results of a comprehensive X-ray spectral analysis of the western hotspot of Pictor A, including new and archival NuSTAR observations totalling 348 ks. We combine this with 278 ks of archival Chandra data from 11 different observations with the Advanced CCD Imaging Spectrometer (ACIS; Garmire et al., 2003) and modeled the combined wide X-ray band spectrum ( keV) of the hot spot for the first time. We also used the maximum likelihood method of Meyer et al. (2023) to test for variability in the NuSTAR and Chandra observations, and discuss the implications of our findings for the X-ray emission origin in Pictor A.
@@ 2.1 NuSTAR observations
The peak of the hot spot emission profile as seen by NuSTAR is slightly offset from the peak observed by Chandra, approximately 4.7” downstream along the jet axis. This offset places the hard X-ray centroid beyond even the radio peak, which is estimated to be downstream of the Chandra position by approximately 1” (Hardcastle et al., 2016). We estimate the uncertainty in the centroid position to be (dx,dy) = (1.72”,2.42”), which we take from the root mean square correction obtained from cross-correlations on the NuSTAR images with the acrosscorr command. However, this offset is most likely due to an asymmetry in the effective PSF combined with the alignment uncertainty. The deconvolved NuSTAR images (see Section C) show a peak consistent with the Chandra images.
@@ 2.2 Chandra observations
The Chandra X-ray Observatory was launched in 1999 as a part of NASA’s Great Observatories program (Weisskopf et al., 2000). The ACIS-S CCD detector (Garmire et al., 2003) onboard has an unrivaled subarcsecond angular resolution in the 0.3 - 10.0 keV energy range, making it ideal for observing kiloparsec-scale extragalactic X-ray jets. Chandra has observed Pictor A in 16 separate pointings, from the initial discovery of the large-scale X-ray jet in 2000 (Wilson et al., 2001) to subsequent observations of the jet and hot spot as recently as 2015 (Hardcastle et al., 2004; Hardcastle & Croston, 2005; Hardcastle et al., 2016). A few observations were excluded from this analysis due to extremely short exposure. Similar to our NuSTAR observations, we grouped observations conducted within week of one another into single epochs. Ultimately, we made use of use 11 Chandra observations which are grouped into 7 epochs and totaling a nominal 330 ks of exposure time. A full list of the Chandra observations used in this analysis, their epoch groupings, PI, project code, observation date, and exposure time before and after filtering can be found in Table 1, alongside the NuSTAR epochs and observations described in the previous section.
All observations were reduced using CIAO version 4.1.2 (Fruscione et al., 2006) and the calibration database CALDB v. 4.9.2. Each observation was first reprocessed using the chandra_repro and filtered to remove all photons outside of the 0.4-8 keV range to minimize background contamination. We use the lc_clean task to bin the good time interval (GTI) data with a bin time of 259.28 s and excise intervals with background flares, which we define as a photon count rate which deviates from the mean count rate by more than 2. We then used this filtered data to produce the 11 sub-pixel images centered on the core of the X-ray jet, one for each epoch, downsizing by a factor of 5 for a pixel scale of 0.0984”. After filtering, the total exposure time is approximately 278 ks.
@@ 2.4 Chandra background
For each Chandra observation, we extracted the background photon counts from a large elliptical region far from the hot spot and bright X-ray core with minor and major axes of 70” and 65” respectively. We used this to generate a background spectral file corresponding to each Chandra observation of the hot spot. Whenever fitting Chandra data in XSPEC, we use the CSTAT statistic which defaults to WSTAT when background data is loaded but not modeled in XSPEC. We bin the spectra such that there are at least 5 counts in the background spectra for each bin to avoid any problematic biases.
@@ 2.5 Spectral modeling of the western hot spot
Powerlaw | powerlaw | N | 2 | Log Unif. | -10 | 0 | |
PhoIndex | 1 | Unif. | 1 | 3 | |||
Broken powerlaw | bknpower | N | 5 | Log Unif. | -10 | 0 | |
PhoIndx1 | 1 | Unif. | 1 | 3 | |||
@@
Double broken powerlaw | bkn2power | N | 6 | Log Unif. | -10 | 0 | |
@@
PhoIndx2 | 3 | Unif. | 0 | 4 | |||
BreakE1 | 2 | Unif. | 5 | ||||
BreakE2 | 2 | Unif. | 5 |

### v2

@@ Abstract
The origin of X-ray emission from the resolved kiloparsec-scale jets and hotspots of many active galactic nuclei (AGN) remains uncertain, particularly where the X-ray emission is separate from the radio-optical synchrotron component. Possible explanations include synchrotron emission from a second electron population and external Compton or synchrotron self-Compton processes – alternatives which imply very different physical conditions within the jet. Until recently, X-ray studies of resolved jets and hotspots have been restricted to below 10 keV, often showing a hard spectral index indicating a spectral peak beyond this energy range. Here we present NuSTAR observations of the nearby powerful radio galaxy Pictor A, in which we clearly detect the western hotspot at approximately 4’ from the host galaxy, the most significant detection of hotspot emission above 10 keV to date. The NuSTAR spectrum is best fit by a single powerlaw of index ; an exponential cut-off gives a 1 lower limit on the cutoff energy of 40.7 keV. We confirm previous findings of variations in the soft X-ray flux detected by Chandra over the 2000 to 2015 period, at a significance of 6.5. This rises to in the common 3-8 keV band using the combined 22-year span of Chandra and NuSTAR observations. The variability of the western Pictor A hotspot strongly confirms the previously argued synchrotron nature of the X-ray emission for the hotspot, while the lower bound to the spectral cutoff energy implies electron energies in the hotspot reach up to at least a few TeV.
@@ 1 Introduction
AGN jets and hotspots are both primarily detected at radio and sometimes IR/optical frequencies where they emit synchrotron radiation from electrons gyrating within the jet magnetic field (e.g., Meisenheimer et al., 1997; Werner et al., 2012). For this reason, jetted AGN are more ‘radio-loud’ compared to the majority of non-jetted AGN. Sources where the jet is aligned in the plane of the sky are known as radio galaxies. Radio galaxies are often divided into two classifications based on jet morphology: Fanaroff Riley type I (FR I; Fanaroff & Riley, 1974) jets tend to be shorter, less collimated, and brightest near the central engine with the jet typically terminating in large, darkened plumes of plasma, while FRII jets are highly collimated and brightest and exhibit terminal hotspots.
@@
One of the major discoveries of Chandra was the detection of bright X-ray emission from resolved radio jets and hotspots on kpc scales (Chartas et al., 2000; Sambruna et al., 2001). In some FRI jets lacking hotspots such as M87 (Harris et al., 2003; Sun et al., 2018) and Cen A (Kraft et al., 2002; Snios et al., 2019), the X-ray emission is roughly consistent with a single spectral component extending from radio to optical to X-rays. In other, often more powerful jets (Harris & Krawczynski, 2006, 2007; Harris et al., 2010; Marshall et al., 2010) the X-ray emission is often ‘anomalous’, meaning that the X-ray flux from the jet and terminal hotspot is too bright and hard-spectrum to be an extension of the primary radio-optical synchrotron component.
For kpc-scale anomalous jets (here we use ’jet’ to mean the resolved components extending up to but not including the hotspot), the initially preferred explanation of inverse Compton scattering of the CMB (IC-CMB) has largely been overturned in favor of a synchrotron origin for the X-rays, at least in jets a lower redshift (e.g. Tavecchio et al., 2000; Celotti et al., 2001; Sambruna et al., 2004; Jester et al., 2006; Meyer et al., 2016; Breiding et al., 2017, 2023; Meyer et al., 2023).
In hotspots the case is somewhat different; thought to represent the jet terminus where the flow finally decelerates to sub-relativistic bulk speed, the IC-CMB contributions are expected to be negligible. While many hotspots are adequately explained as synchrotron self-Compton (SSC) emission (Harris et al., 1994; Hardcastle et al., 2002; Arshakian & Longair, 2000), others are not, requiring magnetic fields grossly out of equipartition (Hardcastle et al., 2004). An SSC origin is also incompatible with the sometimes observed large spatial offsets between the radio and X-rays (Erlund et al., 2007; Reddy et al., 2023). Beaming models based on jet electrons upstream of the termination shock where the flow is still relativistic scattering post-shock radio emission have been proposed (Georganopoulos & Kazanas, 2003; Worrall et al., 2012), but there is limited evidence to support the conclusion that beaming strongly impacts hotspot X-ray emission across the jet population (e.g. Mullin et al., 2008; Hardcastle et al., 2016) and more recent work suggests that the X-ray emission in many hotspots is due synchrotron emission from a second population of electrons, as has been found for many jets (e.g. Mingo et al., 2017).
@@
The source of the kpc-scale X-ray emission for powerful jets and hotspots is a longstanding open question. In some cases alternate models imply orders-of-magnitude differences in jet kinetic power and magnetic field strength. Determining to what extent each possible mechanism contributes across the X-ray jet and hotspot population is thus important also for understanding the energy content of jets and their impact. Given the unsolved origin of the X-ray emission, it is useful to attempt to extend the spectral coverage with observations in the hard X-ray range ( keV). If the emission is synchrotron in origin, locating the turnover point where the spectrum begins to fall can place constraints on the maximum electron energy. This energy range is covered by NuSTAR, but observations of kpc-scale jets are complicated by the fact that NuSTAR lacks the angular resolution (FWHM7.5”) to easily distinguish the kpc-scale jet from the core in most sources. Given these limitations, the outer jet and western hotspot of FR II radio galaxy Pictor A (; 0.69 kpc/”) is nearly the only jet with the sufficient angular separation () from the core to be a viable target for hard X-ray study and the only one bright enough to be detected by NuSTAR in less than approximately 1 Ms of exposure.
@@
Pictor A is among the most well-studied X-ray jets in the literature, and, along with its hotspot, has been the subject of a number of detailed studies over the past two decades (e.g., Wilson et al., 2001; Hardcastle et al., 2004; Hardcastle & Croston, 2005; Migliori et al., 2007; Tingay et al., 2008; Marshall et al., 2010; Hardcastle et al., 2016; Thimmappa et al., 2020). The X-ray jet was first reported by Wilson et al. (2001) and was noted to be X-ray-anomalous by Hardcastle et al. (2004). Despite its bright X-ray emission, it is not a source which we would expect a significant IC-CMB component as it has a low redshift, its (approaching) western jet is not closely inclined to our line-of-sight (), and its counterjet is clearly visible (Hardcastle & Croston, 2005). A beamed inverse-Compton origin for the X-rays from the jet or hotspot would require magnetic field strengths significantly below equipartition values (Hardcastle et al., 2004). The deepest observation of the jet and hotspot to date in the soft X-rays combined 464 ks of Chandra exposure on the western hotspot, finding a spectrum near-equally consistent with a single constant powerlaw spectrum with or a steepening, broken powerlaw spectrum with , , and a break energy at around 2 keV (Hardcastle et al., 2016). One of the motivations for our deep NuSTAR observation was to clarify the presence and location of a possible spectral break or cutoff in the broad X-ray band. Previous research of the hard X-ray emission from the hotspot found no evidence of such a spectral break or cutoff (Sunada et al., 2022); however, as discussed below, the addition of new NuSTAR observations affords us the opportunity to reexamine the spectrum with greater significance.
@@
The jet of Pic A is also noteworthy for the detection of variability in the X-rays. In 2010, Marshall et al. (2010) analyzed three Chandra observations taken in 2000, 2002, and 2009 of 26, 96, and 54 ks of exposure time respectively. They found that the soft X-ray flux of two features located at 34 and 49 kpc away from the core decreased significantly ( and respectively) over the course of the observations. The later Hardcastle et al. (2016) study (with an additional 6 Chandra observations) found the hotspot X-ray flux decreased by over a time period of approximately one month, and confirmed 3 variability in the jet. Recently, Pictor A was included in a large archival study of Chandra jet observations where at least five regions in the jet, including the hotspot, were found to be variable (Meyer et al., 2023); we further expand on the results of that study here using the same methods. The IC-CMB mechanism requires low-energy electrons () with cooling timescales on the order of yrs (Worrall, 2009), so the observed variability in the X-ray flux is clearly incompatible with such an origin. Furthermore, the fast timescales of the variability allow us to probe the physical conditions within the hotspot, most notably the size of the emitting regions (Tingay et al., 2008).
@@
In this paper, we report the results of a comprehensive X-ray spectral analysis of the western hotspot of Pictor A, including new and archival NuSTAR observations totalling 348 ks. We combine this with 313 ks of archival Chandra data from 11 different observations with the Advanced CCD Imaging Spectrometer (ACIS; Garmire et al., 2003) and modeled the combined wide X-ray band spectrum ( keV). A similar research project was undertaken by Sunada et al. (2022); however, access to an additional 300 ks of NuSTAR data allows us to place stricter constraints on the hard and broadband X-ray spectrum than previously possible. We also used the maximum likelihood method of Meyer et al. (2023) to test for variability in the NuSTAR and Chandra observations, and discuss the implications of our findings for the X-ray emission origin in Pictor A.
@@ 2.1 NuSTAR observations
The peak of the hotspot emission profile as seen by NuSTAR at first appears slightly offset from the peak observed by Chandra, by approximately 4-5′′ downstream along the jet axis. If real, this offset would place the hard X-ray centroid beyond even the radio peak, which is estimated to be downstream of the Chandra position by approximately 1′′ (Hardcastle et al., 2016). However, there are two factors which lead us to doubt the reality of the offset: first, there is substantial uncertainty in the centroid position, with dx, dy = 1.72′′, 2.42′′ respectively, where we use the root mean square correction obtained from cross-correlations on the NuSTAR images with the acrosscorr command. In addition, asymmetries in the effective PSF may bias the apparent position of the hotspot – indeed deconvolved NuSTAR images do not show an offset (see further discusion and images in section2.7).
@@ 2.2 Chandra observations
The Chandra X-ray Observatory was launched in 1999 as a part of NASA’s Great Observatories program (Weisskopf et al., 2000). The ACIS-S CCD detector (Garmire et al., 2003) onboard has an unrivaled subarcsecond angular resolution in the 0.3 - 10.0 keV energy range, making it ideal for observing kiloparsec-scale extragalactic X-ray jets. Chandra has observed Pictor A in 16 separate pointings, from the initial discovery of the large-scale X-ray jet in 2000 (Wilson et al., 2001) to subsequent observations of the jet and hotspot as recently as 2015 (Hardcastle et al., 2004; Hardcastle & Croston, 2005; Hardcastle et al., 2016). A few observations were excluded from this analysis due to extremely short exposure. Similar to our NuSTAR observations, we grouped observations conducted within week of one another into single epochs. Ultimately, we made use of use 11 Chandra observations which are grouped into 7 epochs and totaling a nominal 360 ks of exposure time. A full list of the Chandra observations used in this analysis, their epoch groupings, PI, project code, observation date, and exposure time before and after filtering can be found in Table 1, alongside the NuSTAR epochs and observations described in the previous section.
Two observations included in previous deep surveys of Pictor A (Hardcastle et al., 2016; Sunada et al., 2022) have been excluded from this study. The observation corresponding to project code 14357 has the hotspot positioned on a separate chip from the rest of the observation. To minimize complications of differences between the chips, it has been excluded. Similarly, in the observation corresponding to project code 15593, the hotspot is positioned exactly on the chip edge. This complicates the extraction of spectra from the hotspot, and so this observation was also excluded from out analysis.
All observations were reduced using CIAO version 4.1.2 (Fruscione et al., 2006) and the calibration database CALDB v. 4.9.2. Each observation was first reprocessed using the chandra_repro and filtered to remove all photons outside of the 0.4-8 keV range to minimize background contamination. We use the lc_clean task to bin the good time interval (GTI) data with a bin time of 259.28 s and excise intervals with background flares, which we define as a photon count rate which deviates from the mean count rate by more than 2. We then used this filtered data to produce the 11 sub-pixel images centered on the core of the X-ray jet, one for each epoch, downsizing by a factor of 5 for a pixel scale of 0.0984”. After filtering, the total exposure time is approximately 313 ks.
@@ 2.4 Chandra background
For each Chandra observation, we extracted the background photon counts from a large elliptical region far from the hotspot and bright X-ray core with minor and major axes of 70” and 65” respectively. The background region position was identical across all Chandra epochs and always on the same chip as the hotspot. We used this to generate a background spectral file corresponding to each Chandra observation of the hotspot. Whenever fitting Chandra data in XSPEC, we use the CSTAT statistic which defaults to WSTAT when background data is loaded but not modeled in XSPEC. We bin the spectra such that there are at least 5 counts in the background spectra for each bin to avoid any problematic biases.
@@ 2.5 Spectral modeling of the western hotspot
Powerlaw (PL)* | powerlaw | N () | 2 | Log Unif. | -10 | 0 | |
PhoIndex () | 1 | Unif. | 1 | 3 | |||
Broken powerlaw (Bkn. PL) | bknpower | N | 4 | Log Unif. | -10 | 0 | |
PhoIndx1 () | 1 | Unif. | 1 | 3 | |||
BreakE () | 2 | Unif. | |||||
PhoIndx2 () | 3 | Unif. | 1 | 3 | |||
@@
Double broken powerlaw (Dbl. Bkn. PL) | bkn2power | N | 6 | Log Unif. | -10 | 0 | |
@@
BreakE1 () | 2 | Unif. | 5 | ||||
PhoIndx2 () | 3 | Unif. | 0 | 4 | |||
BreakE2 () | 4 | Unif. | 5 | ||||
PhoIndx3 () | 5 | Unif. | 1 | 3 |